OpenStudy (anonymous):

F(2x) can be expressed in the form a+bx. find the values of a and b

OpenStudy (katielong):

someone put this to me if f(x)=1/6 than f(2x)=2(1/6)=2/6=1/3

OpenStudy (anonymous):

did they

OpenStudy (anonymous):

yh but is a and b